OpenStudy (colcaps):

Given: f(x) = 2x2 – 3x + 4 g(x) = x + 3 Find: (f+g)(x)

OpenStudy (anonymous):

(f+g)x=f(x)+g(x)

OpenStudy (anonymous):

2x^2-2x+7

OpenStudy (colcaps):

So do i replace it with the given and solve it all in?

OpenStudy (anonymous):

yup..

OpenStudy (anonymous):

add both

OpenStudy (colcaps):

alright let me try it, i'll solve.

OpenStudy (anonymous):

got it?

OpenStudy (colcaps):

Wait, i got confused in the second equation from the first equation

OpenStudy (colcaps):

2x^2 is 4x right?

OpenStudy (anonymous):

\[2x^2-3x+4+x+3=2x^2-2x+7\]

OpenStudy (anonymous):

no no..

OpenStudy (colcaps):

combine like terms?

OpenStudy (anonymous):

there is no need to "SOLVE" this

OpenStudy (anonymous):

i thought you were talking about combining

OpenStudy (colcaps):

so ill just subtract and add?

OpenStudy (colcaps):

pardon me, im not good with math

OpenStudy (anonymous):

ya..

OpenStudy (anonymous):

jus like numbers

OpenStudy (colcaps):

so lets go back to our equation,2x^2-2x+7 so what will i do first sine it has "x" in it?

OpenStudy (colcaps):

*since

OpenStudy (colcaps):

I mean the 2x^2

OpenStudy (anonymous):

no need to do anything..We wont have to find the x

OpenStudy (colcaps):

2-2 +7?

OpenStudy (anonymous):

what we wrote is the answer

OpenStudy (colcaps):

OH!

OpenStudy (anonymous):

got it?

OpenStudy (colcaps):

do you still need to solve it when you distribute the given to the question?

OpenStudy (anonymous):

Actually we dont solve it unless we are told to find (f+g)(3),(f+g)(4) and stuff...then we substitute that value to our x's

OpenStudy (anonymous):

suppose you are told to find (f+g)(3) then the answer will be 2(3^2)-3(3)+7 =18-9+7=16

OpenStudy (colcaps):

isnt it suppose to be written like this? (2x^2-3x+4+x+3) And what is "(x)"?

OpenStudy (anonymous):

yeah...-3x+x=-2x and 4+3=7..so we wrote it simply like our answer

OpenStudy (anonymous):

you dont have to know what is x

OpenStudy (colcaps):

Waht happened to 2x^2? wow im really confused right now.

OpenStudy (anonymous):

it is there..our final answer is \[2x^2-2x+7\]

OpenStudy (anonymous):

FINAL!!!

OpenStudy (colcaps):

Alright, lets go back because we are having a quiz on functions, ill try to answer this alright?

OpenStudy (colcaps):

Given: f(x) = 2x2 – 3x + 4 g(x) = x + 3 Find: (f+g)(x)

OpenStudy (anonymous):

ok..nothing special happened here..we just added both functions

OpenStudy (colcaps):

So ill distribute the given to my equation right?

OpenStudy (anonymous):

distribute?what do you mean?

OpenStudy (colcaps):

I mean replace it or something

OpenStudy (anonymous):

no..actually we had our final answer..now u dont need to do anything

OpenStudy (colcaps):

Hahaha alright, thanks for helping, ill try to review the steps ey'

OpenStudy (anonymous):

are you thinking of substituting a value to x?

OpenStudy (colcaps):

No, i mean the given will be written in the "(f+g)(x)" problem

OpenStudy (anonymous):

we found the (f+g)(x)

OpenStudy (colcaps):

so (f+g)(x) is = to (2x^2-3x+4+x+3) right?

OpenStudy (colcaps):

also, whats the "(x)"?

OpenStudy (anonymous):

oh (x)?

OpenStudy (colcaps):

yeah

OpenStudy (anonymous):

(x) is just the symbol..it has no value

OpenStudy (anonymous):

you can give any value to it

OpenStudy (colcaps):

so what will i do with it?

OpenStudy (anonymous):

(f+g)(x) is just like writing (f)(x)

OpenStudy (anonymous):

(f+g)and (x) are not seperate things

OpenStudy (colcaps):

Alright, last question, what is 2x^2-3x+4+x+3... So its =4x-3x+4+x+3 =1x+4+x+3 Am i doing it right?

OpenStudy (colcaps):

if yes, should i continue?

OpenStudy (anonymous):

2x^2 is not 4x

OpenStudy (colcaps):

why?

OpenStudy (anonymous):

2x^2=2x^2

OpenStudy (anonymous):

4 is 2^2

OpenStudy (anonymous):

it is 2*(x^2)

OpenStudy (anonymous):

not 2^2 * (x)

OpenStudy (colcaps):

You seperated x^2?

OpenStudy (anonymous):

ya..2 and x^2 are different

OpenStudy (anonymous):

sorry i need to go..bye

OpenStudy (colcaps):

2x^2-3x+4+x+3 =2(x^2)-3x+4+3=x is that right?

OpenStudy (colcaps):

Alright, thanks

OpenStudy (anonymous):

@Colcaps

OpenStudy (anonymous):

you got 2x^2-3x+4+x+3?

OpenStudy (colcaps):

Yes

OpenStudy (anonymous):

2x^2-3x+4+x+3 =2(x^2)-3x+4+3=x ??? it should be this..am i right @Krishnadas 2x^2-2x+7..

OpenStudy (anonymous):

ya you are right @pengembara_bumi1 ...

OpenStudy (anonymous):

so @Colcaps now add -3x and x

OpenStudy (anonymous):

ouh,thanks..hehe ^_^

OpenStudy (colcaps):

-2x?

OpenStudy (anonymous):

ya..now add 4 and 3

OpenStudy (anonymous):

@Colcaps

OpenStudy (anonymous):

7

OpenStudy (colcaps):

4+3 is 7 yeah

OpenStudy (colcaps):

@Krishnadas

OpenStudy (anonymous):

so what are the terms left,2x^2 ,-2x and +7

OpenStudy (anonymous):

adding them gives our final answer

OpenStudy (colcaps):

2-2+7=x?

OpenStudy (anonymous):

no 2x^2-2x+7

OpenStudy (anonymous):

which equals (f+g)(x)

OpenStudy (anonymous):

got it?

OpenStudy (colcaps):

OH!

OpenStudy (anonymous):

why is there a doubt?:D

OpenStudy (anonymous):

okay??

OpenStudy (colcaps):

I thought youll look for the result of it so i got really confused

OpenStudy (colcaps):

Thank you for teaching me mate, that means alot

OpenStudy (anonymous):

this is our result..2x^2-2x+7

OpenStudy (anonymous):

Happy to help:D

OpenStudy (colcaps):

i have another question but its harder than this, ill make a new one

OpenStudy (anonymous):

ok